OpenStudy (anonymous):

1) a=b 2) ab=a^2 3) ab-b^2=a^2-b^2 4) b(a-b)=(a+b)(a-b) 5) b= a+b Reminder the first step where b = 2b So, 1=2. In this case in my opinion the wrong step is that in the third, because you can´t subtract -b^2 from both sides.

Parth (parthkohli):

lol why not?

Parth (parthkohli):

Transition from step 4-5 is wrong. Guess why!

OpenStudy (anonymous):

Tell me ParthKohli.

Parth (parthkohli):

Division by zero. Period.

OpenStudy (anonymous):

Write step-by-step. plz.

Parth (parthkohli):

\[\rm b(a - b) = (a + b)(a - b)\]Since (a - b) = 0,\[\rm 0b = 0(a + b)\]

Parth (parthkohli):

But what this thing shows is that it's divided both sides by 0.

OpenStudy (anonymous):

thanks